OpenStudy (anonymous):

Please help!!! By completing the square in x^2-2x+3, find the minimum value of f(x)=x^2-2x+3

hero (hero):

Do you really need to complete the square just to find the minimum value?

OpenStudy (anonymous):

yeah .

hero (hero):

Why not just simply use the vertex formula?

OpenStudy (anonymous):

I'm not finding the vertex.

hero (hero):

You're finding the minimum value of the function. To do that, use the formula for x coordinate of the vertex: \[x = -\frac{b}{2a}\]

hero (hero):

Then plug that value for x into the function, then find y coordinate.

OpenStudy (anonymous):

can you show me, I don't understand,

hero (hero):

\[x = -\frac{-2}{2(1)} = 1 \\y = f(x) \\y = x^2 -2x + 3 \\y = (1)^2 - 2(1) + 3 \\y = 1 - 2 + 3 \\ y = 3 + 1 - 2 \\ y = 4 - 2 \\y = 2\]

hero (hero):

Thus the minimum point is (1,2)

OpenStudy (anonymous):

Ohhhhh!! Thank Youu!!

hero (hero):

I hope you understand how to do these now.

OpenStudy (zehanz):

Completing the square: f(x) = x^2-2x+3 = (x-1)^2-1+3 = (x-1)^2+2. You can now see that this expression is minimal at x=1. The minimum is f(1) = 2.

hero (hero):

I was just showing her how to do it the easy way first.